# paper=a4paper,margin=2.0cm

### v1

@@ 3.1 Collision between $(1,1,1)$-non-degenerate solitons
As we have mentioned earlier, to study the collision among two ($1,1,1$)-non-degenerate solitons, it is necessary to choose $k_{jI}=l_{jI}$, for $j=1,2$, in the two-soliton solution, which can be deduced from Eqs. (25a)-(25d) by restricting to $N=2$. Then, we also further consider the parametric choices, $k_{jI}=l_{jI}>0$, $k_{jR},l_{jR}>0$, $j=1,2$, and $k_{1I}<k_{2I}$ (or $l_{1I}<l_{2I}$). Note that the singularity condition: $k_{jI}=l_{jI}>0$, for negative nonlinearity coefficients, enforces the ($1,1,1$)-non-degenerate solitons to propagate in the $+x$-direction so that an overtaking collision only occurs in the same direction. Thus, head-on collision is not possible between the ($1,1,1$)-non-degenerate solitons in the case of negative nonlinearity. To deduce the asymptotic forms of these solitons in the long time regimes, we incorporate the asymptotic behaviour of the wave variables, $\eta_{jR}=k_{jR}(x-2k_{jI}t)$ and $\xi_{jR}=l_{jR}(x-2k_{jI}t)$, $j=1,2$, in the corresponding two-soliton solution. By doing so, we have deduced the following asymptotic forms of ($1,1,1$)-non-degenerate solitons.
@@
$e^{i\theta_{1}^{1-}}=\frac{(k_{1}-k_{2})(k_{1}+k_{2})^{\frac{1}{2}}(k_{1}^{*}+k_{2})(k_{1}-l_{2})^{\frac{1}{2}}(k_{1}^{*}-k_{2})^{\frac{1}{2}}(k_{1}^{*}+l_{2})^{\frac{1}{2}}\sqrt{\alpha_{1}^{(1)}}}{(k_{1}^{*}-k_{2}^{*})(k_{1}^{*}+k_{2}^{*})^{\frac{1}{2}}(k_{1}+k_{2}^{*})(k_{1}^{*}-l_{2}^{*})^{\frac{1}{2}}(k_{1}-k_{2}^{*})^{\frac{1}{2}}(k_{1}+l_{2}^{*})^{\frac{1}{2}}\sqrt{\alpha_{1}^{(1)*}}}$, $c_{1}=\frac{1}{2}\log\frac{(k_{1}^{*}-l_{1}^{*})}{(l_{1}-k_{1})}$, $c_{2}=\frac{1}{2}\log\frac{(k_{1}-l_{1})(k_{1}^{*}+l_{1})}{(l_{1}-k_{1})(k_{1}+l_{1}^{*})}$, $e^{i\theta_{2}^{1-}}=\frac{(l_{1}-l_{2})(l_{1}+l_{2})^{\frac{1}{2}}(l_{1}^{*}+l_{2})(k_{2}-l_{1})^{\frac{1}{2}}(l_{1}^{*}-l_{2})^{\frac{1}{2}}(k_{2}+l_{1}^{*})^{\frac{1}{2}}\sqrt{\alpha_{1}^{(2)}}}{(l_{1}^{*}-l_{2}^{*})(l_{1}^{*}+l_{2}^{*})^{\frac{1}{2}}(l_{1}+l_{2}^{*})(k_{2}^{*}-l_{1}^{*})^{\frac{1}{2}}(l_{1}-l_{2}^{*})^{\frac{1}{2}}(k_{2}^{*}+l_{1})^{\frac{1}{2}}\sqrt{\alpha_{1}^{(2)*}}}$, $c_{3}=\frac{1}{2}\log\frac{(k_{1}^{*}-l_{1}^{*})(k_{1}+l_{1}^{*})}{(k_{1}^{*}+l_{1})(k_{1}-l_{1})}$, $b_{1}=(\frac{k_{1R}-l_{1R}}{k_{1R}+l_{1R}})^{\frac{1}{2}}$ $c_{4}=\frac{1}{2}\log\frac{(k_{1}^{*}-l_{1}^{*})(k_{1}^{*}+l_{1})}{(k_{1}+l_{1}^{*})(l_{1}-k_{1})}$, and $\phi_{2}^{-}=\frac{1}{2}\ln\frac{-i(k_{1}-l_{1})|k_{1}-l_{2}|^{2}|k_{1}-k_{2}|^{4}|k_{1}+k_{2}|^{2}|\alpha_{1}^{(1)}|^{2}}{2(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}|k_{1}+l_{2}^{*}|^{2}(k_{1}^{*}+l_{1})|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}}$. In $e^{i\theta_{1,2}^{1-}}$, superscript ($1-$) represents soliton $1$ ($\mathsf{S}_{1}$) before collision and subscripts $(1,2)$ denote the two short-wave components $S^{(1)}$ and $S^{(2)}$, respectively. Note: To get the exact form of ($1,1,1$)-non-degenerate solitons one has to impose the condition: $k_{jI}=l_{jI}$, $j=1,2$, in the constants defined above and the ones appearing in the following asymptotic expressions.
@@
In the above, $a_{21}=\frac{(k_{2}^{*}-l_{2}^{*})^{\frac{1}{2}}}{(k_{2}^{*}+l_{2})^{\frac{1}{2}}}$, $a_{22}=\frac{(k_{2}^{*}-l_{2}^{*})^{\frac{1}{2}}}{(k_{2}+l_{2}^{*})^{\frac{1}{2}}}$, $e^{i\theta_{1}^{2-}}=(\frac{\alpha_{2}^{(1)}}{\alpha_{2}^{(1)*}})^{\frac{1}{2}}$, $e^{i\theta_{2}^{2-}}=(\frac{\alpha_{2}^{(2)}}{\alpha_{2}^{(2)*}})^{\frac{1}{2}}$, $b_{2}=\frac{(k_{2R}-l_{2R})^{\frac{1}{2}}}{(k_{2R}+l_{2R})^{\frac{1}{2}}}$, $\varphi_{1}^{-}=\frac{1}{2}\ln\frac{i(k_{2}-l_{2})|\alpha_{2}^{(2)}|^{2}}{2(k_{2}+l_{2}^{*})(l_{2}-l_{2}^{*})(l_{2}+l_{2}^{*})^{2}}$, $\varphi_{2}^{-}=\frac{1}{2}\ln\frac{-i(k_{2}-l_{2})|\alpha_{2}^{(1)}|^{2}}{2(k_{2}^{*}+l_{2})(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{2}}$, $d_{1}=\frac{1}{2}\log\frac{(k_{2}^{*}-l_{2}^{*})}{(k_{2}-l_{2})}$, $d_{2}=\frac{1}{2}\log\frac{(k_{2}^{*}+l_{2})}{(k_{2}+l_{2}^{*})}$, $d_{3}=\frac{1}{2}\log\frac{(k_{2}^{*}-l_{2}^{*})(k_{2}+l_{2}^{*})}{(k_{2}^{*}+l_{2})(k_{2}-l_{2})}$ and $d_{4}=\frac{1}{2}\log\frac{(k_{2}^{*}-l_{2}^{*})(k_{2}^{*}+l_{2})}{(k_{2}+l_{2}^{*})(k_{2}-l_{2})}$. Here, superscript ($2-$) appears in $e^{i\theta_{1,2}^{2-}}$ refers to soliton 2 ($\mathsf{S}_{2}$) before collision.
@@
$\phi_{1}^{+}=\phi_{1}^{-}-\psi_{1},~{}\phi_{2}^{+}=\phi_{2}^{-}-\psi_{2},~{}\varphi_{1}^{+}=\varphi_{1}^{-}+\Psi_{1},~{}\varphi_{2}^{+}=\varphi_{2}^{-}+\Psi_{2},$ (15a)
@@
$\Delta\Phi_{2}$ $=$ $-\log\frac{|k_{2}-l_{1}||l_{1}-l_{2}|^{2}|l_{1}+l_{2}||k_{1}-l_{2}||k_{1}-k_{2}|^{2}|k_{1}+k_{2}|}{|k_{2}+l_{1}^{*}||l_{1}+l_{2}^{*}|^{2}|l_{1}-l_{2}^{*}||k_{1}+l_{2}^{*}||k_{1}+k_{2}^{*}|^{2}|k_{1}-k_{2}^{*}|}=-\Delta\Phi_{1}.~{}~{}~{}~{}~{}~{}~{}$ (16b)
@@ 3.1.1 Elastic collision property of $(1,1,1)$ non-degenerate solitons
We note that the collision behavior of the $(1,1,1)$ non-degenerate solitons in the system (1) with positive nonlinearity coefficients $\sigma_{l}>0$ is similar to the case discussed above for the negative nonlinearity ($\sigma_{l}<0$) case, except that the collision occurs in the $-x$-direction. In fact, this scenario has already been analyzed in our earlier work [20]. For a mixed choice of nonlinearity coefficients, the LSRI system (1) admits singular higher-order soliton solutions, similar to the case of the one-soliton solution. As a result, analyzing such higher-order solutions is physically meaningless, and therefore, this possibility is excluded from our analysis.
@@ 4 Collision between degenerate and non-degenerate solitons
×$\hskip 14.22636pt\times[\cos\theta\cosh(\eta_{1R}-\xi_{1R}+\Phi_{4}^{-})+i\sin\theta\sinh(\eta_{1R}-\xi_{1R}+\Phi_{4}^{-})]+2(k_{1R}-l_{1R})^{2}\nu_{13}$
×$\hskip 14.22636pt\times\cosh(\eta_{1R}+\xi_{1R}+\Phi_{8}^{-})e^{\frac{\eta_{1R}+\xi_{1R}}{2}}[\cosh\Phi_{5}^{-}\cos\theta+i\sinh\Phi_{5}^{-}\sin\theta]+\nu_{14}\bigg{)},$ (19c)
@@
where $\theta=(k_{1R}^{2}-l_{1R}^{2})t$. The constants such as $d_{ij}$’s, $\nu_{11}$, $\nu_{12}$, $\nu_{13}$, $\nu_{14}$, $\Phi_{j}^{-}$, $i=1,2$, $j=1,2,...,9$, appearing above are defined in Appendix B.
@@
$S^{(l)}\simeq\begin{pmatrix}A_{1}^{-}\\ A_{2}^{-}\end{pmatrix}2k_{2R}\sqrt{k_{2I}}e^{i(\eta_{2I}-\frac{\pi}{2})}\mbox{sech}(\eta_{2R}+\phi^{-}),~{}L\simeq 2k_{2R}^{2}\mbox{ sech}^{2}(\eta_{2R}+\phi^{-}),~{}l=1,2.~{}~{}$ (20)
where $A_{l}^{-}=\alpha_{2}^{(l)}/(\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|\alpha_{2}^{(2)}|^{2})^{1/2}$, $l=1,2$, $\phi^{-}=\frac{1}{2}\ln\frac{(\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|\alpha_{2}^{(2)}|^{2})}{2i(k_{2}+k_{2}^{*})^{2}(k_{2}-k_{2}^{*})}$. Here, the subscript $l$ in $A_{l}^{-}$ denotes the SW components and superscript (--) represents before collision.
@@
$S^{(1,2)}\simeq\begin{pmatrix}A_{1}^{+}\\ A_{2}^{+}\end{pmatrix}2k_{2R}\sqrt{k_{2I}}e^{i(\eta_{2I}+\frac{\pi}{2})}\mbox{sech}(\eta_{2R}+\phi^{+}),~{}L\simeq 2k_{2R}^{2}\mbox{ sech}^{2}(\eta_{2R}+\phi^{+}).~{}~{}$ (22a)
@@
$\chi=(|k_{2}-l_{1}|^{2}|k_{1}+k_{2}^{*}|^{2}|k_{2}+l_{1}|^{2}|k_{1}-k_{2}^{*}|^{2})/(|k_{1}-k_{2}|^{2}|k_{2}+l_{1}^{*}|^{2}|k_{1}+k_{2}|^{2}|k_{2}-l_{1}^{*}|^{2})$, $\phi^{+}=\frac{1}{2}\ln\frac{|k_{1}-k_{2}|^{2}|k_{2}-l_{1}|^{2}\Lambda_{3}}{2i(k_{2}-k_{2}^{*})(k_{2}+k_{2}*)^{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}}$, where $\Lambda_{3}$ is defined in Appendix B. Here, $l$ in $A_{l}^{+}$, $l=1,2$, refers to the SW components and the superscript ($+$) denotes after collision.
@@
From the above analysis, we find that the degenerate soliton $\mathsf{S}_{2}$ strongly influences the structure of ($1,1,1$)-non-degenerate soliton as it is demonstrated in Figs. 7 and 8. As a result, the geometrical structure of non-degenerate soliton $\mathsf{S}_{1}$ gets varied to a different profile structure. For instance, we display such a typical energy sharing collision between a single-humped degenerate soliton and an asymmetric breathing ($1,1,1$)-non-degenerate soliton in Fig. 7. As we can see from this figure, the degenerate soliton $\mathsf{S}_{2}$ undergoes the standard energy sharing collision among the two SW components while colliding with a non-degenerate soliton. That is, the intensity of soliton $\mathsf{S}_{2}$ is suppressed in the SW component $S^{(1)}$ whereas it is enhanced in the other SW component $S^{(2)}$. To preserve the energy conservation, the intensity of soliton $\mathsf{S}_{1}$ is enhanced apprecibly accompanied by a profile transition from a breathing asymmetric shape to a single-humped form with a standard phase shift. In the LW component, the degenerate soliton retains its structural profile, as observed in the purely degenerate case [21], whereas the ($1,1,1$)-non-degenerate soliton $\mathsf{S}_{1}$ exhibits a distinct shape-changing behavior. This novel shape changing collision has not been reported before in the literature for the system (1) with a negative nonlinearity coefficient. During this interesting enegy sharing collision, the polarization constants corresponding to the degenerate soliton of SW components before collision $A_{l}^{-}=\alpha_{2}^{(l)}/(\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|\alpha_{2}^{(2)}|^{2})^{1/2}$, $l=1,2$, change into $A_{1}^{+}=\alpha_{2}^{(1)}/(\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}\chi|\alpha_{2}^{(2)}|^{2})^{1/2}$, $A_{2}^{+}=\alpha_{2}^{(2)}/(\sigma_{1}|\alpha_{2}^{(1)}|^{2}\chi^{-1}+\sigma_{2}|\alpha_{2}^{(2)}|^{2})^{1/2}$, where $\chi=(|k_{2}-l_{1}|^{2}|k_{1}+k_{2}^{*}|^{2}|k_{2}+l_{1}|^{2}|k_{1}-k_{2}^{*}|^{2})/(|k_{1}-k_{2}|^{2}|k_{2}+l_{1}^{*}|^{2}|k_{1}+k_{2}|^{2}|k_{2}-l_{1}^{*}|^{2})$.

### v2

@@ 3.1 Collision between $(1,1,1)$-non-degenerate solitons
As we have mentioned earlier, to study the collision among two ($1,1,1$)-non-degenerate solitons, it is necessary to choose $k_{jI}=l_{jI}$, for $j=1,2$, in the two-soliton solution, which can be deduced from Eqs. (25a)-(25d) by restricting to $N=2$. Then, we also further consider the parametric choices, $k_{jI}=l_{jI}>0$, $k_{jR},l_{jR}>0$, $j=1,2$, and $k_{1I}<k_{2I}$ (or $l_{1I}<l_{2I}$). Note that the singularity condition: $k_{jI}=l_{jI}>0$, for negative nonlinearity coefficients, enforces the ($1,1,1$)-non-degenerate solitons to propagate in the +x+x-direction so that an overtaking collision only occurs in the same direction. Thus, head-on collision is not possible between the ($1,1,1$)-non-degenerate solitons in the case of negative nonlinearity. To deduce the asymptotic forms of these solitons in the long time regimes, we incorporate the asymptotic behaviour of the wave variables, $\eta_{jR}=k_{jR}(x-2k_{jI}t)$ and $\xi_{jR}=l_{jR}(x-2k_{jI}t)$, $j=1,2$, in the corresponding two-soliton solution. By doing so, we have deduced the following asymptotic forms of ($1,1,1$)-non-degenerate solitons.
@@
$e^{i\theta_{1}^{1-}}=\frac{(k_{1}-k_{2})(k_{1}+k_{2})^{\frac{1}{2}}(k_{1}^{*}+k_{2})(k_{1}-l_{2})^{\frac{1}{2}}(k_{1}^{*}-k_{2})^{\frac{1}{2}}(k_{1}^{*}+l_{2})^{\frac{1}{2}}\sqrt{\alpha_{1}^{(1)}}}{(k_{1}^{*}-k_{2}^{*})(k_{1}^{*}+k_{2}^{*})^{\frac{1}{2}}(k_{1}+k_{2}^{*})(k_{1}^{*}-l_{2}^{*})^{\frac{1}{2}}(k_{1}-k_{2}^{*})^{\frac{1}{2}}(k_{1}+l_{2}^{*})^{\frac{1}{2}}\sqrt{\alpha_{1}^{(1)*}}}$, $c_{1}=\frac{1}{2}\log\frac{(k_{1}^{*}-l_{1}^{*})}{(l_{1}-k_{1})}$, $c_{2}=\frac{1}{2}\log\frac{(k_{1}-l_{1})(k_{1}^{*}+l_{1})}{(l_{1}-k_{1})(k_{1}+l_{1}^{*})}$, $e^{i\theta_{2}^{1-}}=\frac{(l_{1}-l_{2})(l_{1}+l_{2})^{\frac{1}{2}}(l_{1}^{*}+l_{2})(k_{2}-l_{1})^{\frac{1}{2}}(l_{1}^{*}-l_{2})^{\frac{1}{2}}(k_{2}+l_{1}^{*})^{\frac{1}{2}}\sqrt{\alpha_{1}^{(2)}}}{(l_{1}^{*}-l_{2}^{*})(l_{1}^{*}+l_{2}^{*})^{\frac{1}{2}}(l_{1}+l_{2}^{*})(k_{2}^{*}-l_{1}^{*})^{\frac{1}{2}}(l_{1}-l_{2}^{*})^{\frac{1}{2}}(k_{2}^{*}+l_{1})^{\frac{1}{2}}\sqrt{\alpha_{1}^{(2)*}}}$, $c_{3}=\frac{1}{2}\log\frac{(k_{1}^{*}-l_{1}^{*})(k_{1}+l_{1}^{*})}{(k_{1}^{*}+l_{1})(k_{1}-l_{1})}$, $b_{1}=(\frac{k_{1R}-l_{1R}}{k_{1R}+l_{1R}})^{\frac{1}{2}}$ $c_{4}=\frac{1}{2}\log\frac{(k_{1}^{*}-l_{1}^{*})(k_{1}^{*}+l_{1})}{(k_{1}+l_{1}^{*})(l_{1}-k_{1})}$, and $\phi_{2}^{-}=\frac{1}{2}\ln\frac{-i(k_{1}-l_{1})|k_{1}-l_{2}|^{2}|k_{1}-k_{2}|^{4}|k_{1}+k_{2}|^{2}|\alpha_{1}^{(1)}|^{2}}{2(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}|k_{1}+l_{2}^{*}|^{2}(k_{1}^{*}+l_{1})|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}}$. In $e^{i\theta_{1,2}^{1-}}$, superscript (11-) represents soliton 11 ($\mathsf{S}_{1}$) before collision and subscripts $(1,2)$ denote the two short-wave components $S^{(1)}$ and $S^{(2)}$, respectively. Note: To get the exact form of ($1,1,1$)-non-degenerate solitons one has to impose the condition: $k_{jI}=l_{jI}$, $j=1,2$, in the constants defined above and the ones appearing in the following asymptotic expressions.
@@
In the above, $a_{21}=\frac{(k_{2}^{*}-l_{2}^{*})^{\frac{1}{2}}}{(k_{2}^{*}+l_{2})^{\frac{1}{2}}}$, $a_{22}=\frac{(k_{2}^{*}-l_{2}^{*})^{\frac{1}{2}}}{(k_{2}+l_{2}^{*})^{\frac{1}{2}}}$, $e^{i\theta_{1}^{2-}}=(\frac{\alpha_{2}^{(1)}}{\alpha_{2}^{(1)*}})^{\frac{1}{2}}$, $e^{i\theta_{2}^{2-}}=(\frac{\alpha_{2}^{(2)}}{\alpha_{2}^{(2)*}})^{\frac{1}{2}}$, $b_{2}=\frac{(k_{2R}-l_{2R})^{\frac{1}{2}}}{(k_{2R}+l_{2R})^{\frac{1}{2}}}$, $\varphi_{1}^{-}=\frac{1}{2}\ln\frac{i(k_{2}-l_{2})|\alpha_{2}^{(2)}|^{2}}{2(k_{2}+l_{2}^{*})(l_{2}-l_{2}^{*})(l_{2}+l_{2}^{*})^{2}}$, $\varphi_{2}^{-}=\frac{1}{2}\ln\frac{-i(k_{2}-l_{2})|\alpha_{2}^{(1)}|^{2}}{2(k_{2}^{*}+l_{2})(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{2}}$, $d_{1}=\frac{1}{2}\log\frac{(k_{2}^{*}-l_{2}^{*})}{(k_{2}-l_{2})}$, $d_{2}=\frac{1}{2}\log\frac{(k_{2}^{*}+l_{2})}{(k_{2}+l_{2}^{*})}$, $d_{3}=\frac{1}{2}\log\frac{(k_{2}^{*}-l_{2}^{*})(k_{2}+l_{2}^{*})}{(k_{2}^{*}+l_{2})(k_{2}-l_{2})}$ and $d_{4}=\frac{1}{2}\log\frac{(k_{2}^{*}-l_{2}^{*})(k_{2}^{*}+l_{2})}{(k_{2}+l_{2}^{*})(k_{2}-l_{2})}$. Here, superscript (22-) appears in $e^{i\theta_{1,2}^{2-}}$ refers to soliton 2 ($\mathsf{S}_{2}$) before collision.
@@
$\hskip-71.13188pt\phi_{1}^{+}=\phi_{1}^{-}-\psi_{1},~\phi_{2}^{+}=\phi_{2}^{-}-\psi_{2},~\varphi_{1}^{+}=\varphi_{1}^{-}+\Psi_{1},~\varphi_{2}^{+}=\varphi_{2}^{-}+\Psi_{2},$ (15a)
@@
$\hskip-42.67912pt\Delta\Phi_{2}$ $=$ $-\log\frac{|k_{2}-l_{1}||l_{1}-l_{2}|^{2}|l_{1}+l_{2}||k_{1}-l_{2}||k_{1}-k_{2}|^{2}|k_{1}+k_{2}|}{|k_{2}+l_{1}^{*}||l_{1}+l_{2}^{*}|^{2}|l_{1}-l_{2}^{*}||k_{1}+l_{2}^{*}||k_{1}+k_{2}^{*}|^{2}|k_{1}-k_{2}^{*}|}=-\Delta\Phi_{1}.$~~~~~~~ (16b)
@@ 3.1.1 Elastic collision property of $(1,1,1)$ non-degenerate solitons
We note that the collision behavior of the $(1,1,1)$ non-degenerate solitons in the system (1) with positive nonlinearity coefficients $\sigma_{l}>0$ is similar to the case discussed above for the negative nonlinearity ($\sigma_{l}<0$) case, except that the collision occurs in the x-x-direction. In fact, this scenario has already been analyzed in our earlier work [20]. For a mixed choice of nonlinearity coefficients, the LSRI system (1) admits singular higher-order soliton solutions, similar to the case of the one-soliton solution. As a result, analyzing such higher-order solutions is physically meaningless, and therefore, this possibility is excluded from our analysis.
@@ 4 Collision between degenerate and non-degenerate solitons
$\hskip 14.22636pt\times[\cos\theta\cosh(\eta_{1R}-\xi_{1R}+\Phi_{4}^{-})+i\sin\theta\sinh(\eta_{1R}-\xi_{1R}+\Phi_{4}^{-})]+2(k_{1R}-l_{1R})^{2}\nu_{13}$
$\hskip 14.22636pt\times\cosh(\eta_{1R}+\xi_{1R}+\Phi_{8}^{-})e^{\frac{\eta_{1R}+\xi_{1R}}{2}}[\cosh\Phi_{5}^{-}\cos\theta+i\sinh\Phi_{5}^{-}\sin\theta]+\nu_{14}\bigg{)},$ (19c)
@@
where $\theta=(k_{1R}^{2}-l_{1R}^{2})t$. The constants such as $d_{ij}$’s, $\nu_{11}$, $\nu_{12}$, $\nu_{13}$, $\nu_{14}$, $\Phi_{j}^{-}$, $i=1,2$, j=1,2,,9j=1,2,...,9, appearing above are defined in Appendix B.
@@
$\hskip-14.22636ptS^{(l)}\simeq\begin{pmatrix}A_{1}^{-}\\ A_{2}^{-}\end{pmatrix}2k_{2R}\sqrt{k_{2I}}e^{i(\eta_{2I}-\frac{\pi}{2})}\mbox{ sech}(\eta_{2R}+\phi^{-}),~L\simeq 2k_{2R}^{2}\mbox{ sech}^{2}(\eta_{2R}+\phi^{-}),~l=1,2.$~~ (20)
where $A_{l}^{-}=\alpha_{2}^{(l)}/(\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|\alpha_{2}^{(2)}|^{2})^{1/2}$, $l=1,2$, $\phi^{-}=\frac{1}{2}\ln\frac{(\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|\alpha_{2}^{(2)}|^{2})}{2i(k_{2}+k_{2}^{*})^{2}(k_{2}-k_{2}^{*})}$. Here, the subscript $l$ in $A_{l}^{-}$ denotes the SW components and superscript (-) represents before collision.
@@
$\hskip-14.22636ptS^{(1,2)}\simeq\begin{pmatrix}A_{1}^{+}\\ A_{2}^{+}\end{pmatrix}2k_{2R}\sqrt{k_{2I}}e^{i(\eta_{2I}+\frac{\pi}{2})}\mbox{ sech}(\eta_{2R}+\phi^{+}),~L\simeq 2k_{2R}^{2}\mbox{ sech}^{2}(\eta_{2R}+\phi^{+}).$~~ (22a)
@@
$\chi=(|k_{2}-l_{1}|^{2}|k_{1}+k_{2}^{*}|^{2}|k_{2}+l_{1}|^{2}|k_{1}-k_{2}^{*}|^{2})/(|k_{1}-k_{2}|^{2}|k_{2}+l_{1}^{*}|^{2}|k_{1}+k_{2}|^{2}|k_{2}-l_{1}^{*}|^{2})$, $\phi^{+}=\frac{1}{2}\ln\frac{|k_{1}-k_{2}|^{2}|k_{2}-l_{1}|^{2}\Lambda_{3}}{2i(k_{2}-k_{2}^{*})(k_{2}+k_{2}*)^{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}}$, where $\Lambda_{3}$ is defined in Appendix B. Here, $l$ in $A_{l}^{+}$, $l=1,2$, refers to the SW components and the superscript (++) denotes after collision.
@@
From the above analysis, we find that the degenerate soliton $\mathsf{S}_{2}$ strongly influences the structure of ($1,1,1$)-non-degenerate soliton as it is demonstrated in Figs. 8 and 9. As a result, the geometrical structure of non-degenerate soliton $\mathsf{S}_{1}$ gets varied to a different profile structure. For instance, we display such a typical energy sharing collision between a single-humped degenerate soliton and an asymmetric breathing ($1,1,1$)-non-degenerate soliton in Fig. 8. As we can see from this figure, the degenerate soliton $\mathsf{S}_{2}$ undergoes the standard energy sharing collision among the two SW components while colliding with a non-degenerate soliton. That is, the intensity of soliton $\mathsf{S}_{2}$ is suppressed in the SW component $S^{(1)}$ whereas it is enhanced in the other SW component $S^{(2)}$. To preserve the energy conservation, the intensity of soliton $\mathsf{S}_{1}$ is enhanced apprecibly accompanied by a profile transition from a breathing asymmetric shape to a single-humped form with a standard phase shift. In the LW component, the degenerate soliton retains its structural profile, as observed in the purely degenerate case [21], whereas the ($1,1,1$)-non-degenerate soliton $\mathsf{S}_{1}$ exhibits a distinct shape-changing behavior. This novel shape changing collision has not been reported before in the literature for the system (1) with a negative nonlinearity coefficient. During this interesting enegy sharing collision, the polarization constants corresponding to the degenerate soliton of SW components before collision $A_{l}^{-}=\alpha_{2}^{(l)}/(\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|\alpha_{2}^{(2)}|^{2})^{1/2}$, $l=1,2$, change into $A_{1}^{+}=\alpha_{2}^{(1)}/(\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}\chi|\alpha_{2}^{(2)}|^{2})^{1/2}$, $A_{2}^{+}=\alpha_{2}^{(2)}/(\sigma_{1}|\alpha_{2}^{(1)}|^{2}\chi^{-1}+\sigma_{2}|\alpha_{2}^{(2)}|^{2})^{1/2}$, where $\chi=(|k_{2}-l_{1}|^{2}|k_{1}+k_{2}^{*}|^{2}|k_{2}+l_{1}|^{2}|k_{1}-k_{2}^{*}|^{2})/(|k_{1}-k_{2}|^{2}|k_{2}+l_{1}^{*}|^{2}|k_{1}+k_{2}|^{2}|k_{2}-l_{1}^{*}|^{2})$.